OpenStudy (anonymous):

Which linear function represents the line given by the point-slope equation y + 1 = –3(x – 5)? f(x) = –3x – 6 f(x) = –3x – 4 f(x) = –3x + 16 f(x) = –3x + 14

Mehek (mehek14):

first use distribution \(-3(x-5)\\-3*x=-3x\\-3*-5=15\\y+1=-3x+15\)

Mehek (mehek14):

what is 15 - 1 @jonestyrese15

OpenStudy (anonymous):

i think the answer is d

Mehek (mehek14):

you are correct ^_^

OpenStudy (imtiaz7):

f(x)= -3x+14

Mehek (mehek14):

@Imtiaz7 sort of late :P

OpenStudy (anonymous):

can u help me with one more

Mehek (mehek14):

sure

OpenStudy (anonymous):

The equation of a linear function in point-slope form is y – y1 = m(x – x1). Harold correctly wrote the equation y = 3(x – 7) using a point and the slope. Which point did Harold use? When Harold wrote his equation, the point he used was (7, 3). When Harold wrote his equation, the point he used was (0, 7). When Harold wrote his equation, the point he used was (7, 0). When Harold wrote his equation, the point he used was (3, 7).

OpenStudy (imtiaz7):

Yup :P@Mehek14

Mehek (mehek14):

you have to find which point is on the line y = 3(x – 7)

Mehek (mehek14):

which is (7,0)

OpenStudy (anonymous):

so c

Mehek (mehek14):

yes

OpenStudy (anonymous):

Which equations and/or functions represent the graphed line? Check all that apply. f(x) = 0.2x - 4 f(x) = 0.5x + 2 f(x) = x + 2 y – 3 = (x – 2) y – 1 = 0.5(x + 2)

Mehek (mehek14):

f(x) = 0.5x + 2

OpenStudy (anonymous):

thts the only one

Mehek (mehek14):

y – 1 = 0.5(x + 2)

Mehek (mehek14):

those 2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Which linear function represents the line given by the point-slope equation y + 1 = –3(x – 5)? f(x) = –3x – 6 f(x) = –3x – 4 f(x) = –3x + 16 f(x) = –3x + 14